OpenStudy (anonymous):

Calculate the pressure exerted on the floor when an elephant who weighs 6750 N stands on one foot which has an area of 25 inches squared.

OpenStudy (ybarrap):

Pressure = Force/Area = 6,750 N/25 \(in^2\times 12~ in^2/1ft^2\times 1 ft^2/0.0929~m^2\\=\cfrac{6,750\times 12}{0.0929}~Pascals\)

OpenStudy (anonymous):

thanks soooo much!

OpenStudy (ybarrap):

yw